### v1

# AnyBimanual: Transferring Unimanual Policy for General Bimanual Manipulation

Abstract
footnotetext: : equal contribution. : corresponding author.
Performing general language-conditioned bimanual manipulation tasks is of great importance for many applications ranging from household service to industrial assembly. However, collecting bimanual manipulation data is expensive due to the high-dimensional action space, which poses challenges for conventional methods to handle general bimanual manipulation tasks. In contrast, unimanual policy has recently demonstrated impressive generalizability across a wide range of tasks because of scaled model parameters and training data, which can provide sharable manipulation knowledge for bimanual systems. To this end, we propose a plug-and-play method named AnyBimanual, which transfers pre-trained unimanual policy to general bimanual manipulation policy with few bimanual demonstrations. Specifically, we first introduce a skill manager to dynamically schedule the skill representations discovered from pre-trained unimanual policy for bimanual manipulation tasks, which linearly combines skill primitives with task-oriented compensation to represent the bimanual manipulation instruction. To mitigate the observation discrepancy between unimanual and bimanual systems, we present a visual aligner to generate soft masks for visual embedding of the workspace, which aims to align visual input of unimanual policy model for each arm with those during pretraining stage. AnyBimanual shows superiority on 12 simulated tasks from RLBench2 with a sizable 12.67% improvement in success rate over previous methods. Experiments on 9 real-world tasks further verify its practicality with an average success rate of 84.62%.
Figure 1: AnyBimanual enables plug-and-play transferring from pre-trained unimanual policies to bimanual manipulation policy, which preserves the generalizability with the proposed skill scheduling framework.
1 Introduction
Bimanual systems play an important role in robotic manipulation due to the high capacity of completing diverse tasks in household service [72], robotic surgery [33], and component assembly in factories [9]. Compared to unimanual systems, bimanual systems expand the workspace and are able to handle more complex manipulation tasks by stabilizing the target with one arm and interacting with that using another arm [30, 45]. Even for tasks that unimanual policies can handle, bimanual systems are often more efficient because multiple action steps can be simultaneously accomplished [31]. Since modern robotic applications require the robot to interact with different tasks and objects, it is desirable to design a generalizable policy model for bimanual manipulation.
To enhance the generalization ability of the manipulation agent, prior works present to leverage the high-level reasoning and semantic understanding capabilities of foundation models like Large Language Models (LLMs) and Vision Language Models (VLMs) to breakdown tasks into executable sub-tasks that can be solved by external low-level controllers [34, 26, 44, 37, 27], which thus struggles with contact-rich tasks that require complex and precise low-level motion. To generalize to contact-rich tasks, recent methods [41, 51, 1] tend to learn robotic foundation models directly from large-scale teleoperation data [40, 17, 55], which has shown impressive generalizability across various unimanual tasks. However, bimanual demonstrations are extremely expensive to acquire in real-world, which often require specialized teleoperation systems with additional sensors and fine-grained calibration with high human labor cost [73, 22, 18, 12, 56, 62]. To address this challenge, recent methods aim to simplify the learning budget by exploiting human inductive bias such as parameterized atomic movements that detail the position and rotation [15, 5] or assigning stabilizing and acting roles for each arm [45, 54, 24], thereby reducing the need for extensive expert data. Nevertheless, shareable atomic movements and fixed cooperation patterns struggle to generalize across different bimanual manipulation tasks, which limits the deployment scenario of these classes of methods.
In this paper, we propose a plug-and-play module named AnyBimanual that transfers any pre-trained unimanual policy to bimanual manipulation policy with limited demonstrations. Since unimanual policy model [50, 39] has demonstrated impressive generalization ability across tasks due to the large model sizes and numerous training demonstrations, we realize high generalizability across diverse language-conditioned bimanual manipulation tasks by mining and transferring the commonsense knowledge in pre-trained unimanual policies. More specifically, we first introduce a skill manager that dynamically schedules discovered skill representations. Skill representations are formed by skill primitives that store shareable manipulation knowledge across embodiments, with task-oriented importance weights and compensation. To enhance the transferability of the pre-trained unimanual policy in bimanual manipulation tasks, the observation discrepancy between unimanual and bimanual systems should be minimized. We propose a voxel aligner to generate spatial soft masks to highlight relevant visual clues for different arms, whose goal is to align the visual input of the unimanual policy model for each arm with those during the pretraining stage. We evaluate AnyBimanual on a comprehensive task suite composed of $12$ simulated tasks from RLBench2 [31] and $9$ real-world tasks, where our method surpasses the previous state-of-the-art method by a large margin. The contributions are summarized as follows:
  • We propose a model-agnostic plug-and-play framework named AnyBimanual that transfers an arbitrary pre-trained unimanual policy to generalizable bimanual manipulation policy with limited bimanual demonstrations.
  • We introduce a skill manager to dynamically schedule skill representations for unimanual policy transferring and a visual aligner to mitigate the observation discrepancy between unimanual and bimanual systems for transferability enhancement.
  • We conduct extensive experiments of $12$ tasks from RLBench2 and $9$ tasks from the real world. The results demonstrate that our method achieves a higher success rate than the state-of-the-art methods.
2 Related Work
Generalizable Bimanual Manipulation. Bimanual manipulation agents [6, 38, 22, 71, 68, 27, 45, 24, 31, 11] are able to handle a large variety of tasks by predicting a trajectory of bimanual operation, which is of great significance in complex applications from household service [72], robotic surgery [33], to component assembly in factories [9]. To achieve multi-task learning for generalizable Bimanual manipulation, earlier studies attempted to leverage the emerged general understanding and reasoning capacities of pre-trained foundation models like LLMs [53] and VLMs [10], where the foundation model was prompted to generate a high-level plan for low-level executors. However, the performance of directly leveraging foundation models in a training-free manner is bottlenecked by the predefined low-level executor, which struggles to generalize to more contact-rich tasks like straightening a rope. To overcome this challenge, robotic foundation models [41, 46, 17, 8, 7, 51] that pre-trained on large-scale real-world demonstrations were proposed under the unimanual setting, which has shown high generalizability across everyday manipulation tasks. However, bimanual tasks demand precise coordination of two high degree-of-freedom arms, making the teleoperation of demonstrations for training generalizable policies also costly. Although some recent approaches [73, 22, 18, 16, 65, 13] have developed more specialized teleoperation systems to reduce these costs, scaling up demonstrations for high generalization ability remains a challenge. To address the limited availability of demonstrations, alternative methods [60, 45, 30] proposed to simplify the learning of bimanual policies by decoupling the bimanual system into a stabilizing arm and an acting arm. Nevertheless, these methods often rely on predefined roles for each arm, which precludes their applicability to tasks requiring more flexible collaboration patterns. In contrast to these approaches, our work presents a novel method that transfers generalizable unimanual policies to bimanual tasks, which eliminates the necessity for explicit inductive bias like role specification.
Figure 2: The overall pipeline of AnyBimanual, which primarily consists of a skill manager and a perception manager. The skill manager adaptively coordinates primitive skills for each robot arm, while the perception manager mitigates the distributional shift from unimanual to bimanual by decomposing the 3D voxel observation for each arm.
Skill-based Methods. Skill learning [70] is the process where intelligent agents acquire new abilities that are transferable across different tasks, which is of great significance for cross-task generalization. Thus, skill learning is being attractive in enhancing the generalizability of different models, such as game agents [57], robotic manipulation [43], and autonomous driving [20]. The initial attempt to utilize skill learning was orchestrating a set of predefined skill [48], which hindered their scalability to unseen tasks. To overcome this limitation, [43, 15] proposed to learn shareable skill primitives from data. For example, skill diffuser [43] introduced a hierarchical planning framework that integrates learnable skill embedding into conditional trajectory generation, which realized accurate execution of diverse compositional tasks. In the field of bimanual manipulation, skill learning was dominated by handcrafted primitives. For example, [4, 19, 29, 3, 23, 2, 66, 21, 14, 64, 32] propose to utilize parameterized atomic movements to shrink the high dimensionality of the bimanual action space, which has shown impressive performance on templated bimanual manipulation tasks. While the predefined atomic movements did boost the success rate on specific tasks, they are often difficult for even human users to specify, which restricts the deployment scenarios of these methods. In this paper, we propose leveraging learnable skill primitives to represent the learned commonsense of the pre-trained unimanual policy, so that the knowledge can be transferred across different levels of tasks.
3 Approach
In this section, we first introduce required preliminaries (Section 3.1), and then we present a pipeline overview (Section 3.2). Then, we introduce a skill manager (Section 3.3) that schedules skills to each arm to form general bimanual manipulation policies. We build a visual aligner (Section 3.4) to mitigate the observation discrepancy between bimanual and unimanual systems for better generalization. Finally, we outline the training objectives (Section 3.5).
3.1 Problem Formulation
The task of policy learning for bimanual manipulation can be defined as follows. To complete a wide range of manipulation tasks specified in natural language, the bimanual agent is required to interactively predict the actions of both end-effectors based on the visual observation and robot states, where the motion is acquired by a low-level planner (e.g., RRT-Connect). The observation $o_{t}$ at the $t_{th}$ time step includes the voxel $v_{t}$ converted from RGB and depth images [31, 50] and the robot proprioception $p_{t}$. The action $a_{t}$ for each end-effector at the $t_{th}$ time step contains the location $a_{\text{trans}}$, orientation $a_{\text{rot}}$, gripper open state $a_{\text{open}}$ and usage of collision avoidance in the motion planner $a_{\text{col}}$. For the training data, human demonstrators produce a limited set of $M$ offline expert trajectories $\mathcal{D}=\{(o_{1},a^{\text{left}}_{1},a^{\text{right}}_{1}),...,(o_{M},a^{\text{left}}_{M},a^{\text{right}}_{M})\}$ for each task instruction $l$, where $a^{\text{arm}}_{t},\text{arm}\in\mathcal{A}=\{\text{left},\text{right}\}$ demonstrates the actions for left and right grippers. Existing methods directly learn the policy model from expert demonstrations, which have shown effectiveness in single-task settings. However, due to the high cost of data collection in bimanual systems, the scarcity of expert demonstrations limits the generalizability of these methods across tasks. To address this, we present to transfer pre-trained generalizable unimanual policy for general bimanual manipulation.
3.2 Overall Pipeline
Overall pipeline of our AnyBimanual method is shown in Figure 2. For the language branch, we employed a pre-trained text encoder [49] to parse the bimanual instruction to language embeddings with high-level semantics, where the skill manager schedules the skill primitives with composition and compensation to enhance the language embeddings that instruct relevant subtasks for different arms. Therefore, the pre-trained unimanual policy model can be prompted to generate feasible manipulation policy for each arm with high generalization ability across tasks with the sharable manipulation knowledge. For the visual branch, we voxelize the RGB-D input to the voxel space as observation, and a 3D sparse voxel encoder is utilized to tokenize the voxel observation for informative volumetric representation. The visual aligner generates a soft spatial mask to align the visual representation of unimanual policy model with its representation during pretraining, so that the observation discrepancy between unimanual and bimanual systems can be minimized for policy transferability enhancement. We employ two pre-trained unimanual models to predict the left and right robot actions based on the text embeddings and visual representations, where the pre-trained unimanual policy can be multi-modal transformer-based policy [50, 41, 8, 7, 17] or diffusion-based policy [39, 51].
3.3 Scheduling Unimanual Skill Primitives
In order to transfer unimanual manipulation policy to bimanual settings without generalizability drops, we propose a skill manager to decompose the action policies from unimanual foundation models into skill primitives and integrate primitives for bimanual systems. However, the given offline expert demonstrations $\mathcal{D}$ do not contain any explicit intermediate skill primitives or sub-task boundaries, but only low-level end-effector poses and high-level natural language instruction are provided. Therefore, we design an automatic skill discovery method in an unsupervised manner to learn skill representations and their schema from offline bimanual manipulation datasets during training. In the test phase, the skill manager predicts different weighted combinations of primitive skills to orchestrate each arm given high-level language instruction, which enables effective transfer of pre-trained unimanual policy to diverse bimanual manipulation tasks.
Skill Manager. We start with a discrete primitive skill set $\mathcal{Z}=\{z_{1},z_{2},...,z_{K}\}$, where $K$ is a hyper-parameter that indicates the number of skill primitives. To realize end-to-end skill discovery and scheduling, each potential skill is an implicit embedding $z_{k}\in\mathbb{R}^{D}$, which can be initialized with the corresponding language template tokens of the pre-trained unimanual policy to mitigate the domain gap.
Figure 3: Shareable skills across unimanual and bimanual settings. We observe that bimanual tasks are often originated from the combination of unimanual sub-tasks, which thus can be solved by effectively coordinating unimanual skills synchronously or asynchronously.
By combining the primitives from the skill set, the language embedding for the unimanual policy model can be represented as a linear combination of these primitives. Hence, the reconstructed language embedding as skill representation can be expressed as:
$\hat{l}^{\text{left}}_{t}=\sum^{K}_{k=1}\hat{w}^{\text{left}}_{k,t}z_{k}+\epsilon^{\text{left}}_{t},\quad\hat{l}^{\text{right}}_{t}=\sum^{K}_{k=1}\hat{w}^{\text{right}}_{k,t}z_{k}+\epsilon^{\text{right}}_{t}$ (1)
where $\hat{l}^{\text{arm}}_{t}$ is the decomposed unimanual language embedding for one arm of the bimanual system, and $\hat{w}^{\text{arm}}_{t}\in\mathbb{R}^{K}$ denotes the importance weight for linear combination for both values of $\text{arm}\in\mathcal{A}$. We also introduce a task-oriented compensation $\epsilon^{\text{arm}}_{t}\in\mathbb{R}^{D}$ to introduce the embodiment-specific knowledge for the policy transfer. The upscript arm of the variables can be selected from left or right to indicate the embodiment in the bimanual systems. As depicted in Figure 3, considering a bimanual task Handover, it can be explicitly solved by scheduling two unimanual primitive skills, i.e., the left arm Place the block to the right gripper while the right arm Pick it from the left gripper.
Though every language embedding that passed through the pre-trained single policy can be represented as a linear combination of the skill set, the combination weights that specify the importance of each skill are dynamic across the task completion process. We propose to parametrize a multi-model transformer named skill manager to dynamically predict the combination weight for each arm at each time step. Therefore, our skill manager $f$ can be formulated as $(\hat{w}^{\text{left}}_{t},\epsilon^{\text{left}}_{t},\hat{w}^{\text{right}}_{t},\epsilon^{\text{right}}_{t})=f_{\theta}(v_{t},l,p_{t})$, which takes the overall bimanual visual and language embeddings, proprioception as input, and assigns the reconstructed unimanual language embedding for each arm as output to schedule the skill primitive of each arm dynamically. $\theta$ represents the learnable parameters. Finally, the combined skill primitives are concatenated with the initial bimanual language embedding to enhance the global context, which is then forwarded to the corresponding unimanual policy.
Learning Generalizable Skill Representations. To update the skill library, we expect the discorverd skill representations are informative to encode fundamental robot motions that are sharable across a variety of tasks, thereby enhancing the generalizability of our framework. To realize this, the learning objective of the skill manager is designed as a sparse representation problem [61]:
$\mathcal{L}_{\text{skill}}=\|\hat{w}^{\text{left}}\|_{1}+\|\hat{w}^{\text{right}}\|_{1}+\lambda_{\epsilon}(\|\epsilon^{\text{left}}\|_{2,1}+\|\epsilon^{\text{right}}\|_{2,1})$ (2)
where $\|\cdot\|_{1}$, and $\|\cdot\|_{2,1}$ denote the $\ell_{1}$ and $\ell_{2,1}$ norm, respectively. $\lambda_{\epsilon}$ is a hyper-parameter that balances the denoising term. By minimizing this sparse regularization term, the skill manager is encouraged to reconstruct the skill representation by selecting fewer skill primitives, which is further surrogated by minimizing a differentiable $\ell_{1}$ regularization [52]. Therefore, the skill subspaces are required to be orthogonal and disjoint with each other to reconstruct the language embedding with sparse combination and compensation, which implicitly enforces each skill representation to capture an independent fundamental motion.
3.4 Aligning Unimanual Visual Representations
Visual Aligner. Despite the skill manager enabling generalization in the language modality, the distributional shift from the unimanual to bimanual workspace in terms of the visual context still may harm the model performance. To mitigate the observation discrepancy, we present a visual aligner $q$ that predict two spatial soft masks at each step $t$ to edit the voxel space so that the decomposed subspace of unimanual policy model for each arm aligns with those during pretraining stage: $(\hat{v}^{\text{left}}_{t},\hat{v}^{\text{right}}_{t})=q_{\theta}(v_{t},l,p_{t})$. The decomposed observation represents the locality of the workspace, which is then augmented by the bimanual observation to form the final visual embedding:
$v^{\text{left}}_{t}=(\hat{v}^{\text{left}}_{t}\odot v_{t})\oplus v_{t},\quad v^{\text{right}}_{t}=(\hat{v}^{\text{right}}_{t}\odot v_{t})\oplus v_{t},$ (3)
where $\odot$ is the element-wise multiplication, and $\oplus$ refers to the concatenation operator. As a result, the augmented visual representations for each arm contains both embodiment-specific information and the global context, which is then passed through the two unimanual policy models to decode the optimal bimanual action.
Learning Aligned Visual Representations. Our goal is to mitigate the visual domain gap between the unimanual and bimanual setting, so that the pre-trained commonsense knowledge in unimanual policy can be transferred with high adaptation ability. Since we can not access the unimanual pretraining data in common usages, we instead impose a mutually exclusive prior to the visual aligner. This prior is regularized by optimizing a Jensen-Shannon (JS) divergence regularization term:
$\mathcal{L}_{\text{voxel}}=-D_{KL}(\hat{v}^{\text{left}}_{t}\|\hat{v}^{\text{right}}_{t})/2-D_{KL}(\hat{v}^{\text{right}}_{t}\|\hat{v}^{\text{left}}_{t})/2$ (4)
where $D_{KL}$ means the Kullback-Leibler (KL) divergence operator. To provide further explanation, bimanual manipulation tasks often involve asynchronous collaboration that requires the left and right arm to attend to different areas of the whole workspace to act as different roles, such as stabilizing and acting. As a result, the mutually exclusive division of the entire bimanual workspace will naturally separate one arm and its target from the other, which highly resembles the unimanual configuration. Hence by maximizing the divergence between the two soft masks, the voxel input of the bimanual manipulation agent can be disentangled into unimanual visual representations that align with those in the pretraining phase effectively.
3.5 Learning Objectives
The decomposed skill and volumetric representation are used to pass through the two pre-trained unimanual policies to predict the optimal actions of the two end-effectors. We assume access to a pre-trained unimanual policy $p$, which is fundamentally a multi-model multi-task neural network that takes visual and language embedding as inputs and outputs end-effector actions. Our AnyBimanual is a model-agnostic plug-and-play method, which indicates that the architecture of pre-trained unimanual policy $p$ is flexible in different architectures such as multi-modal transformer-based policies [50, 41] and diffusion policies [39, 51]. To supervise the model with the provided expert demonstrations for behavior cloning, we leverage the cross-entropy loss to learn accurate action prediction for each arm:
$\mathcal{L}_{\text{BC}}=\sum_{\text{arm}\in\mathcal{A}}CE(p^{\text{arm}}_{\text{trans}},p^{\text{arm}}_{\text{rot}},p^{\text{arm}}_{\text{open}},p^{\text{arm}}_{\text{col}})$ (5)
where $p^{\text{arm}}_{\text{trans}},p^{\text{arm}}_{\text{rot}},p^{\text{arm}}_{\text{open}},p^{\text{arm}}_{\text{col}}$ represents the distribution of the ground-truth actions for translation, rotation, gripper openness, and collision avoidance for the corresponding robot arm, respectively. The behavior cloning loss is then combined with the two regularization terms described above to learn informative skill manager and visual aligner. To sum up, the training objective of AnyBimanual is:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{BC}}+\lambda_{\text{skill}}\mathcal{L}_{\text{skill}}+\lambda_{\text{voxel}}\mathcal{L}_{\text{voxel}},$ (6)
where $\lambda_{\text{skill}}$ and $\lambda_{\text{voxel}}$ refer to hyper-parameters that balance the importance of the regularizations.
Table 1: Multi-Task Test Results in Simulator. Average success rates (%) of general bimanual manipulation agents trained with $20$ or $100$ demonstrations per task and evaluated over $25$ episodes. ‘LF’ means the leader-follower architecture that transfers unimanual manipulation policy for bimanual manipulation. AnyBimanual enables plug-and-play transfers for multiple unimanual manipulation policies.
pick
laptop
pick
plate
straighten
rope
lift
ball
lift
tray
push
box
[…]
RVT [28]-LF 0 0 0 0 0 0 4 12 0 4 12 16
RVT-LF + AnyBimanual 0 0 4 8 0 8 4 16 0 8 20 24
PerAct [50]-LF 0 0 0 4 4 8 4 12 0 4 8 16
PerAct-LF + AnyBimanual 0 4 0 4 4 4 8 12 0 8 12 24
PerAct2 [31] 0 8 4 16 0 4 8 40 0 4 0 0
PerAct2 + Pretraining 0 4 0 20 4 12 16 12 0 12 0 8
PerAct + AnyBimanual 8 8 12 32 8 24 4 32 4 12 28 16
put in
fridge
press
buttons
handover
item
sweep to
dustpan
take out
tray
handover
easy
[…]
RVT [28]-LF 0 0 12 12 0 0 0 0 0 4 0 0
RVT-LF + AnyBimanual 0 0 16 20 0 4 0 4 4 4 0 8
PerAct [50]-LF 0 0 0 8 0 0 8 28 0 0 0 4
PerAct-LF + AnyBimanual 0 4 4 8 0 0 16 32 0 0 0 8
PerAct2 [31] 0 0 4 12 8 0 8 0 0 0 28 24
PerAct2 + Pretraining 0 0 44 64 0 4 0 8 8 4 16 28
PerAct + AnyBimanual 0 8 48 84 8 8 4 4 8 4 24 28
Table 2: Comparison of AnyBimanual with Different Techniques. We categorize the long-term tasks that require more than $6.5$ keyframes to Long, tasks that involve multiple variations to Generalized and tasks that involve synchronization of both arms to Sync for further interpretability.
Row ID Skill Manager Visual Aligner Long Generalized Sync Average
1 - - 6.86 6.00 12.00 9.00
2 10.85 32.00 8.00 14.33
3 12.57 44.00 14.00 18.67
4 13.71 42.00 13.00 19.00
5 14.28 46.00 21.00 21.67
4 Experiments
In this section, we first introduce the experimental setups (Section 4.1). Then, we transfer various unimanual methods to bimanual manipulation via AnyBimanual, and compare them with the state-of-the-art approaches to show the superiority (Section 4.2). We conduct an ablation study to evaluate the validity of the proposed components (Section 4.3). We further interpret the learned skill representation and decomposed volumetric representation by visualization (Section 4.4). Finally, we report the real-robot results to demonstrate the effectiveness of AnyBimanual in real-world applications (Section 4.5).
4.1 Experiment Setup
Simulation. For benchmarking, our simulation experiments are conducted on RLBench2 [31], a bimanual version extended from the widely-used RLBench [36] benchmark in prior works [35, 69, 47, 28, 63, 39]. Following the setup in [31], we utilize $12$ language-conditioned bimanual manipulation tasks varying from different challenge levels. The diverse task suite requires the agent to acquire and correctly schedule shareable skills to achieve high success rates, rather than merely imitating limited expert demonstrations. For observation, we employ six cameras with a resolution of $256\times 256$ to cover the entire workspace. During the training phase, we provide $20$ or $100$ demonstrations for each task, and we evaluate $25$ episodes per task in the testing set to mitigate the randomness.
Real Robot. The real-world setup for our experiments involves two Universal Robots UR5e manipulators equipped with Robotiq 2F-85 grippers, controlled by two Xbox joysticks for collecting demonstrations. A calibrated front RGB-D Realsense camera provides $640\times 480$ resolution images at $30$ Hz for observation. We collect $30$ real-world human demonstrations per task for training, while the evaluations are conducted using a Nvidia RTX 4080 GPU.
Baselines. We compare our AnyBimanual with the state-of-the-art general bimanual manipulation agents, including the voxel-based method PerAct2 [31] and its leader-follower version PerAct-LF, both are modified from the well-known unimanual policy PerAct [50], as well as the multi-view image-based method RVT [28]-LF. To exclude the influence of model parameters, we also implement a counterpart that directly combines two pre-trained PerAct [50] policies. Note that the proposed method is model-agnostic, which supports different communication mechanisms between single-arm policies, and thus we transfer all $3$ baselines to validate the versatility of AnyBimanual. The evaluation metric is the task success rate, which is defined as the percentage of episodes where the agent successfully completes the instructed goal within $25$ steps.
Implementation Details. Following the common training recipe [50, 31, 28], we use the SE(3) observation augmentation for the expert demonstrations in the training set to improve the robustness of the agents. For fair comparisons, all compared methods are trained for $100$k iterations on two NVIDIA RTX 3090 GPUs with a total batch size of $4$. We use the LAMB optimizer [67] with a constant learning rate of $5\times 10^{-4}$ to update model parameters, in line with the previous arts [31, 50, 28].
Figure 4: Visualization of AnyBimanual. This figure shows in different key timesteps, how the skill manager dynamically schedules skill weights and how the visual aligner decomposes volumetric observation. We use a logarithmic scale for visualization.
Figure 5: Real-World Tasks. The real-world experiments are performed in a tabletop setup with objects randomized in location every episode. AnyBimanual can simultaneously conduct $9$ complex real-world bimanual manipulation tasks with one model. Different colors mean different success rates.
4.2 Comparison with the State-of-the-Art Methods
In this section, we compare our AnyBimanual with previous stat-of-the-art approaches on RLBench tasksuite. Table 1 presents a comparison of the average success rates for each task and the average performance is shown in Figure 1. Our method achieves the highest overall performance, with an average success rate of $21.67$%, setting a new state-of-the-art in general bimanual manipulation. AnyBimanual leverages the knowledge distilled from unimanual models and successfully transfers it to guide general bimanual manipulation. As a result, our method outperforms the cutting-edge bimanual method PerAct2 by a significant improvement of $12.67$% on average. With a limited set of $20$ demonstrations, our method still defeats the baseline with a sizable margin of $8.00$%. Note that we aim to train multi-task agents to handle a wide range of tasks within a single model, hence the performance is lower than the reported in the original paper. Especially, we observe that AnyBimanual shows greater improvement in long-horizon tasks like put in fridge, multi-variations tasks like press buttons, and tasks demand synchronous coordinations like pick plate. Moreover, AnyBimanual enables plug-and-play transferring of PerAct-LF and RVT, which also leads to relative boosts of $25.87$% ($7.00$% to $9.00$%) and $116.65$% ($4.00$% to $8.67$%) on average. However, incorporating AnyBimanual slightly affects the performance on the simple, short-horizon Lift ball, due to the additional complexity in input processing.
Figure 6: Unseen Data Generalization. We include multiple distractors in real-world experiments, and find AnyBimanual generalizes to these settings successfully by unlocking the commonsense held by unimanual base models.
4.3 Ablation Study
Our AnyBimanual framework leverages a skill manager to dynamically coordinate unimanual skills, while mitigating the distributional shift between unimanual and bimanual visual inputs with a visual aligner. We conduct an ablation study in Table 2 by transferring the powerful unimanul baseline PerAct [50] to bimanual tasks. We first implement a vanilla baseline without any of the proposed techniques (Row $1$), where we load the pre-trained PerAct model and directly finetune the baseline to predict bimanual action, which shows enhanced long-horizon task execution and generalizability with an unneglected performance drop in tasks that require proper coordination.
Skill Manager. By employing the skill manager in the experiment, we observe that the average success rate increases by $4.67$% (Row $4$ vs. Row $2$). Notably, in tasks requiring long-term manipulation within the Long category, it significantly outperforms the vanilla version, demonstrating the skill manager’s effectiveness in long-horizon tasks.
Visual Aligner. Additionally, we incorporate the spatial soft masking from the visual aligner, resulting in a substantial performance improvement of 4.34% (Row $3$ vs. Row $2$). Although the inclusion of spatial soft masking mechanism may slightly affect performance in simpler tasks due to increased input processing complexity, it leads to significant gains in overall task success. The gains are even more notable in Generalized and Sync tasks, which demand a high degree of adaptability to scene variations and synchronization for cooperative actions. The results demonstrate the visual aligner preserves the generalizability stored in the unimanual policy perfectly, facilitating seamless coordination between the arms. By integrating all techniques in our AnyBimanual, the success rate improves from $9.00$% to $21.67$% (Row $5$), validating the importance of properly mining unimanual knowledge for bimanual manipulation.
4.4 Qualitative Analysis
[…]
4.5 Real-Robot Results
Table 3: Real-world Results. Average success rates (%) of our multi-task model trained and evaluated on $9$ real-world tasks.
Task Sync. Object Variation Episode Keyframe Average
Lift Cabinet 1 1 5 3.0 100.0
Fold Clothes 1 1 5 3.0 100.0
Pick&Place Sync 7 3 15 3.0 80.0
Handover Bowl 1 1 5 4.0 100.0
Wash Dishes 2 1 5 4.0 100.0
Play Ping Pong 3 1 5 3.0 100.0
Rotate Toothbrush 3 1 5 4.0 20.0
Typewriting 4 1 5 6.0 100.0
Pick&Place Async 5 3 15 4.0 80.0
We further validate our approach through real-robot experiments. We train one multi-task AnyBimanual agent that transfers the unimanual policy PerAct [50] to $9$ real-world tasks and report the average success rates on Table 3. Guided by  [42], these tasks are designed to cover different challenge levels, such as synchronous and asynchronous coordination, short and long-horizon execution, etc. In the real-robot experiments, our keyframes are manually extracted, providing more stable task guidance compared to heuristic extraction methods in simulation. The overall success rate across $65$ test episodes in total is $84.62$%, which illustrates a significant practicality of AnyBimanual in real-world settings. Especially, in multi-variant tasks that require high generalizability like Pick&Place Sync and Pick&Place Async shown in Figure 6, our AnyBimanual completes $80$% tasks of different initial positions, object colors, object size, lightning and backgrounds successfully. The failure cases occur on tasks that demand precise rotation, e.g., Rotate Toothbrush, which could be mitigated by leveraging high-capacity unimanual policy [41] or balancing the weight of the rotation term in behavior cloning.
5 Conclusion
In this paper, we have introduced AnyBimanual, a framework designed to transfer pre-trained unimanual manipulation policies to multi-task bimanual manipulation with few bimanual demonstrations. We develop a skill manager to dynamically schedule skill primitives discovered from unimanual policies, enabling their effective adaptation for bimanual tasks. To address the observation discrepancies between unimanual and bimanual systems, we propose a visual aligner that generates spatial soft masks, aligning the visual embeddings of each arm with those used during the pretraining stage of the unimanual policy model. Extensive experiments across $12$ simulated and $9$ real-world tasks demonstrate the effectiveness of AnyBimanual. The limitations are discussed in the supplementary file.
Supplementary Material
6 Additional Experimental Details
6.1 Simulation
We utilize RLBench2 [31] as our main simulated task suite for mulit-task learning. Table 5 is an overview of the $12$ selected tasks we use in the experiments. Table 4 is an overview of the $18$ selected tasks from RLBench [36] used to pretrain the unimanual checkpoint. The task variations include randomly sampled colors, sizes, counts, placements, and categories of objects. Other properties vary depending on the specific task. Furthermore, objects are randomly arranged on the tabletop within a certain range, adding to the diversity of the tasks. The visual observation includes $6$ RGB-D frames of $256\times 256$ resolutions (i.e., front, shoulder left, shoulder right, wrist left, wrist right and overhead), which are shown in Figure 7. In the ablation study, we group the RLBench2 tasks of Table 5 into $3$ categories according to their key challenges. The task groups include:
  • The Long group includes long-term tasks that requires more than $6.5$ keyframes. The included tasks are: pick plate, pick laptop, put in fridge, handover item, sweep to dustpan, take out of tray and handover easy.
  • The Generalized group includes tasks with multiple variations, which are differentiated by instructions. The included tasks are: press buttons and handover item.
  • The Sync group requires precise coordination, which often causes failures due to asynchronous manipulation. The included tasks are: lift ball, lift tray, straighten rope and push box.
Table 4: Unimanual Tasks. This table shows the $18$ unimanual tasks in RLBench that are used to pretrain the unimanual policy.
Task # of Variations # of Average Keyframes Human Instruction Template
open drawer 3 3.0 “open the   drawer”
slide block 4 4.7 “slide the block to   target”
sweep to dustpan 2 4.6 “sweep dirt to the   dustpan”
meat off grill 2 5.0 “take the   off the grill”
turn tap 2 2.0 “turn   tap”
put in drawer 3 12.0 “put the item in the   drawer”
close jar 20 6.0 “close the   jar”
drag stick 20 6.0 “use the stick to drag the cube onto the   target”
stack blocks 60 14.6 “stack     blocks”
screw bulb 20 7.0 “screw in the   light bulb”
put in safe 3 5.0 “put the money away in the safe on the   shelf”
place wine 3 5.0 “stack the wine bottle to the   of the rack”
put in cupboard 9 5.0 “put the   in the cupboard”
sort shape 5 5.0 “put the   in the shape sorter”
push buttons 50 3.8 “push the   button, [then the   button]”
insert peg 20 5.0 “put the ring on the   spoke”
stack cups 20 10.0 “stack the other cups on top of the   cup”
place cups 3 11.5 “place   cups on the cup holder”
Table 5: Bimanual Tasks. This table shows the $12$ bimanual tasks in RLBench2 that are used to fine-tune the bimanual policy.
Task # of Variations # of Average Keyframes Human Instruction Template
pick laptop 1 7.2 “pick up the notebook.”
pick plate 1 6.6 “pick up the plate.”
straighten rope 1 5.9 “straighten the rope.”
lift ball 1 4.0 “lift the ball.”
lift tray 1 5.1 “lift the tray.”
push box 1 2.1 “push the box to the red area.”
put in fridge 1 7.8 “put the bottle into the fridge.”
press buttons 5 4.0 “push the   and   button.”
handover item 5 7.6 “hand over the   item.”
sweep to dustpan 1 7.3 “sweep the dust to the pan.”
take out tray 1 8.7 “take tray out of oven.”
handover easy 1 7.5 “hand over the item.”
6.2 Real-Robot
Hardware Setup. The real robot setup uses two Universal Robots UR5e manipulators, each equipped with a Robotiq 2F-85 gripper. See Figure 8 for reference. For the perception, we use a Realsense RGB-D camera mounted on a tripod, positioned to mimic the viewpoint of human eyes during bimanual tasks. The Realsense provides RGB-D images at a resolution of 640x480 with a frame rate of 30 Hz. The extrinsics between the camera and right arm base-frame are calibrated using the easy_handeye package [^1] https://github.com/IFL-CAMP/easy_handeye. Additionally, an ARUCO marker [^2] https://github.com/pal-robotics/aruco_ros attached to the UR5e’s end-effector is employed to aid in the calibration process.
Figure 7: Visual Observation in RLBench2. We adopt $6$ RGB-D cameras to cover the whole workspace.
Figure 8: Real-Robot Setup with one RealSense L515 RGB-D Camera and Two UR5e Manipulators.
Tasksuite Descriptions. We provide a detailed description of the real-world tasks we used to evaluate our AnyBimanual in Figure 9, where we follow [42] to classify our $9$ real-world experiments into $5$ categories, covering $5$ bimanual collaboration patterns.
  • Lift Cabinet requires the agent to simultaneously hold two sides of the cabinet to lift it up, which poses challenges to synchronous coordination.
  • Fold Clothes requires the agent to simultaneously grasp and fold the flexible clothes, which involves soft manipulation and synchronous coordination.
  • pick&place Sync requires the agent to simultaneously pick up two colored cubes specified by the human instructions, and then place them in corresponding boxes, which involves semantic understanding.
  • Handover Bowl requires the agent to lift a bowl and hand it over to the other hand, posing challenges in coordinating the handover between hands.
  • Wash Dishes requires the agent to pick up a bowl and then apply detergent to it, testing the asynchronous coordination of both hands.
  • Play Ping Pong requires the agent to pick up both a ball and a racket, toss the ball with one hand and hit it with the racket in the other in a high toss serve, testing asynchronous coordination.
  • Rotate Toothbrush requires the agent to pick up a toothbrush and a cup simultaneously, then rotate the toothbrush 180 degrees to drop it into the cup, challenging the precision of rotation and coordination of both hands.
  • Type Writing requires the agent to sequentially type the letters ”robot” on a keyboard using both hands, testing long-horizon coordination.
  • Pick & Place Async requires the agent to simultaneously pick up two cubes of different colors and then place them into the same bowl, testing the asynchronous coordination of both hands.
Figure 9: Real-Robot Task Taxonomy. Following [42], we classify the reported $9$ real-world tasks into $5$ categories according to their collaboration patterns.
Figure 10: Real-world Keyframes. We manually select keyframes from the collected trajectories to simplify training.
Data Collection and Preprocessing. We collect demonstrations with two Xbox joysticks. Each gamepad is a 6-DoF controller. The gamepads adjust the velocity of the arm’s end-effector to translate and rotate in all directions, with reference to the arm’s base frame. For robot control, we utilize the Universal Robots ROS Driver [^3] https://github.com/UniversalRobots/Universal_Robots_ROS_Driver. After collecting the complete trajectory, we extract keyframes from the trajectory manually to simplify the robot learning. Two examples of the extracted keyframes are shown in Figure 10. The total time consumption of collecting one episode including controlling the robot arms and keyframe extraction is $\sim 1.5$ minutes.
Figure 11: SE(3) Augmentation. We slightly modify the SE(3) augmentation strategy to ensure that the perturbed ground-truth actions of both arms are reasonable.
Data Augmentation. In line with [50, 31], we turn on SE(3) augmentation to boost the robustness of all learned policies. Based on the open source code of [31], we modify the perturbation strategy to ensure that the perturbed ground-truth actions of both arms are reasonable in Figure 11. We plan to release our code for more details.
Policy Deployment. During the testing phase, we first randomize the positions of related objects in the current task. After resetting the manipulators to the home positions, we query the policy with the language instruction, initial observation and proprioception to obtain the next best poses of two end-effectors. We then use Moveit [^4] https://github.com/moveit/moveit to reach the poses and query the policy again with new observations, until task completion or reaching the maximal episode length.
7 Additional Implementation Details
7.1 Hyperparameters
The hyperparameters used in AnyBimanual are shown in Table 6. To ensure that the auxiliary objectives $\mathcal{L}_{\text{skill}}$ and $\mathcal{L}_{\text{voxel}}$ remain in the same magnitude as $\mathcal{L}_{\text{BC}}$, we set $\lambda_{\text{skill}}=0.0001$ and $\lambda_{\text{voxel}}=0.001$. Other hyperparameters are in line with previous works [50, 31] for fair comparisons.
Table 6: A default set of hyper-parameters.
Config Value
Training iteration $100$k
Image size $256\times 256\times 3$
Voxel size $100\times 100\times 100$
Batch size $4$
Optimizer LAMB [67]
Learning rate $0.0005$
Weight decay $0.000001$
$\lambda_{\text{skill}}$ 0.0001
$\lambda_{\text{voxel}}$ 0.001
Number of skill primitives $K$ 18
Dimension of Skill Representations $D$ 512
Figure 12: Skill Clustering. We cluster the skill representations to further interpret the discovered skills of our AnyBimanual.
Figure 13: Comparions of the rollouts of PerAct2 and our AnyBimanual. AnyBimanual demonstrates complex collaboration patterns by incorporating the skill manager and the visual aligner to schedule the manipulators correctly.
Figure 14: Error Modes in Real-world Rollouts. We study the common failure cases of our AnyBimanual in real-world settings, where the incorrect end-effector actions are marked with red circles. We also provide the correct actions with green circles for reference.
7.2 Architectures
Skill Manager. We fuse voxel embeddings and language embeddings by projecting both to a hidden size of $256$ and concatenating them. After layer normalization, the combined sequence is processed by a frozen GPT-2 Transformer configured with a hidden size of $128$ and $4$ attention heads. The transformer’s output is passed through a linear layer to produce logits over 18 classes, and softmaxed to obtain probabilities. These probabilities are then used to reconstruct the skill representation via a weighted sum over the skill primitives.
Visual Aligner. We employ a lightweight module to process input embeddings through a Conv1D-ReLU block, then split into two branches to generate left and right masks via additional Conv1D-ReLU layers. After element-wise multiplication of these masks with the original input, we incorporate residual connections by adding the masked outputs back to the input embeddings.
8 Additional Experimental Results
8.1 Skill Clustering
The goal of the skill manager is to obtain the language embedding via linear skill representations, so that the pre-trained unimanual policy model (i.e., PerAct) can be prompted to generate feasible manipulation actions. In Figure 12, we cluster the skill representations while evaluating the last checkpoint of PerAct+AnyBimanual in RLBench2. We can observe that the learned skill representations are reasonably clustered into $18$ categories, which corresponds to the number of pre-trained unimanual tasks. We also visualize the related unimanual skill of each cluster by finding the nearest unimanual task embedding. The clustering results further interpret that the proposed skill manager enables compact skill representation learning with rich semantics for dynamic scheduling.
8.2 Rollout Comparisons
We present a qualitative example of the generated action sequences in Figure 13, comparing PerAct2 and our AnyBimanual method. In this case, the instruction is “Push the box to the red area”. The PerAct2 agent fails to touch the box and only mimics the expert’s forward-pushing motion, resulting in an incomplete action. In contrast, our AnyBimanual agent ensures both arms make contact with the box and successfully push it into the red area, demonstrating the superior ability of our method in having each arm correctly identify contact with the object and successfully complete the task.
8.3 Error Modes
To further study the limitations and risks of the proposed method, we visualize the common failure cases in Figure 14. Common error modes include coarse control, which is often viewed in Pick&Place Sync and Pick&Place Async. In the upper case, the robot fails to grasp the green block due to a small end-effector deviation, which may be caused by inconsistent data collection. In the middle case, the robot struggles to generalize to different backgrounds and object sizes, which leads to an out-of-distribution movement of the left robot arm. In the bottom case, the low-level motion planner cannot handle the severe rotation predicted by the agent, which can result in a dangerous cup fall. We hope that the analysis can enlighten future works in bimanual policy transferring and prevent potential risks in real-world deployments.
8.4 Real-world Videos
We provide $9$ additional comprehensive real-world episodes generated by our AnyBimanual in the attached video file (demo.mp4). Note that all $9$ tasks are completed by a single model with different natural language instructions to specify the current task.
8.5 Real-world Experiments with Limited Data.
To further validate the effectiveness and practicality of AnyBimanual with limited expert data, we train AnyBimanual on $9$ real-world tasks, each with $5$ demonstrations. As per the main paper, we evaluate each variation over $5$ episodes, and report the success rates of each task. The results are illustrated in Table 7, where AnyBimanual still maintains an acceptable average success rate of $53.33$%.
Table 7: Real-world Results with $5$ Demonstrations Per Task. We train AnyBimanual on $9$ real-world tasks with $5$ demonstrations for each task, and evaluate each variation over $5$ episodes.
      Task       Success Rate (%)
      Lift Cabinet       80.0
      Fold Clothes       80.0
      Pick&Place Sync       80.0
      Handover Bowl       20.0
      Wash Dishes       20.0
      Play Ping Pong       40.0
      Rotate Toothbrush       20.0
      Typewriting       60.0
      Pick&Place Async       80.0
9 Discussions on Limitations
Cross-embodiment Skill Transferring. AnyBimanual needs careful visual alignment between unimanual and bimanual systems to ensure robust policy transfer, which means it is not designed for cross-embodiment skill transferring. Even though the 7-DoF end-effector action space can be used with various robots, differences in the visual appearance of unimanual and bimanual robots may limit generalizability. Because the visual aligner only attempts to align the observation by segmenting out unimanual workspace from the bimanual system, without considering their appearance differences. This can be addressed by visual inpainting [59] or transferring cross-embodiment unimanual policies [58].
Cross-task Generalization. Although AnyBimanual enables general multi-task bimanual manipulation with few demonstrations, it struggles to perform zero-shot generalization to unseen tasks. This is mainly due to the end-to-end skill manager and visual aligner still necessitating expert demonstrations to learn proper coordination patterns and unlock the capacity of unimanual policy. Explicit methods such as leveraging large language models [25] to orchestrate unimanual policies can realize weak zero-shot generalization, but struggle with contact-rich tasks without learning.

### v2

@@
 Abstract
-footnotetext: : equal contribution. : corresponding author.
 Performing general language-conditioned bimanual manipulation tasks is of great importance for many applications ranging from household service to industrial assembly.
@@
 In contrast, unimanual policy has recently demonstrated impressive generalizability across a wide range of tasks because of scaled model parameters and training data, which can provide sharable manipulation knowledge for bimanual systems.
-To this end, we propose a plug-and-play method named AnyBimanual, which transfers pre-trained unimanual policy to general bimanual manipulation policy with few bimanual demonstrations.
-Specifically, we first introduce a skill manager to dynamically schedule the skill representations discovered from pre-trained unimanual policy for bimanual manipulation tasks, which linearly combines skill primitives with task-oriented compensation to represent the bimanual manipulation instruction.
+To this end, we propose a plug-and-play method named AnyBimanual, which transfers pretrained unimanual policy to general bimanual manipulation policy with few bimanual demonstrations.
+Specifically, we first introduce a skill manager to dynamically schedule the skill representations discovered from pretrained unimanual policy for bimanual manipulation tasks, which linearly combines skill primitives with task-oriented compensation to represent the bimanual manipulation instruction.
 To mitigate the observation discrepancy between unimanual and bimanual systems, we present a visual aligner to generate soft masks for visual embedding of the workspace, which aims to align visual input of unimanual policy model for each arm with those during pretraining stage.
-AnyBimanual shows superiority on 12 simulated tasks from RLBench2 with a sizable 12.67% improvement in success rate over previous methods.
+AnyBimanual shows superiority on 12 simulated tasks from RLBench2 with a sizable 17.33% improvement in success rate over previous methods.
 Experiments on 9 real-world tasks further verify its practicality with an average success rate of 84.62%.
-Figure 1: AnyBimanual enables plug-and-play transferring from pre-trained unimanual policies to bimanual manipulation policy, which preserves the generalizability with the proposed skill scheduling framework.
 1 Introduction
 Bimanual systems play an important role in robotic manipulation due to the high capacity of completing diverse tasks in household service [72], robotic surgery [33], and component assembly in factories [9].
-Compared to unimanual systems, bimanual systems expand the workspace and are able to handle more complex manipulation tasks by stabilizing the target with one arm and interacting with that using another arm [30, 45].
-Even for tasks that unimanual policies can handle, bimanual systems are often more efficient because multiple action steps can be simultaneously accomplished [31].
+Compared to unimanual systems, bimanual systems enlarge the workspace and are able to handle more complex manipulation tasks by stabilizing the target with one arm and interacting with that using another arm [30, 45].
+Even for the tasks that unimanual policies can handle, bimanual systems are often more efficient because multiple action steps can be simultaneously accomplished [31].
 Since modern robotic applications require the robot to interact with different tasks and objects, it is desirable to design a generalizable policy model for bimanual manipulation.
+Figure 1: AnyBimanual enables plug-and-play transferring from pretrained unimanual policies to bimanual manipulation policy, which preserves the generalizability with the proposed skill scheduling framework.
 To enhance the generalization ability of the manipulation agent, prior works present to leverage the high-level reasoning and semantic understanding capabilities of foundation models like Large Language Models (LLMs) and Vision Language Models (VLMs) to breakdown tasks into executable sub-tasks that can be solved by external low-level controllers [34, 26, 44, 37, 27], which thus struggles with contact-rich tasks that require complex and precise low-level motion.
 To generalize to contact-rich tasks, recent methods [41, 51, 1] tend to learn robotic foundation models directly from large-scale teleoperation data [40, 17, 55], which has shown impressive generalizability across various unimanual tasks.
-However, bimanual demonstrations are extremely expensive to acquire in real-world, which often require specialized teleoperation systems with additional sensors and fine-grained calibration with high human labor cost [73, 22, 18, 12, 56, 62].
-To address this challenge, recent methods aim to simplify the learning budget by exploiting human inductive bias such as parameterized atomic movements that detail the position and rotation [15, 5] or assigning stabilizing and acting roles for each arm [45, 54, 24], thereby reducing the need for extensive expert data.
+However, bimanual demonstrations are extremely expensive to acquire in real-world, which often need specialized teleoperation systems with additional sensors and fine-grained calibration with high human laborer cost [73, 22, 18, 12, 56, 62].
+To address this challenge, recent methods aim to simplify the learning budget by exploiting human inductive bias like parameterized atomic movements that detail the position and rotation [15, 5] or assigning stabilizing and acting roles for each arm [45, 54, 24], thereby reducing the need for extensive expert data.
 Nevertheless, shareable atomic movements and fixed cooperation patterns struggle to generalize across different bimanual manipulation tasks, which limits the deployment scenario of these classes of methods.
 In this paper, we propose a plug-and-play module named AnyBimanual that transfers any pre-trained unimanual policy to bimanual manipulation policy with limited demonstrations.
-Since unimanual policy model [50, 39] has demonstrated impressive generalization ability across tasks due to the large model sizes and numerous training demonstrations, we realize high generalizability across diverse language-conditioned bimanual manipulation tasks by mining and transferring the commonsense knowledge in pre-trained unimanual policies.
+Since unimanual policy model [50, 39] has demonstrated impressive generalization ability across tasks due to the large model sizes and numerous training demonstrations, we realize high generalizability across diverse language-conditioned bimanual manipulation tasks by mining and transferring the commonsense knowledge in pretrained unimanual policies.
 More specifically, we first introduce a skill manager that dynamically schedules discovered skill representations.
 Skill representations are formed by skill primitives that store shareable manipulation knowledge across embodiments, with task-oriented importance weights and compensation.
-To enhance the transferability of the pre-trained unimanual policy in bimanual manipulation tasks, the observation discrepancy between unimanual and bimanual systems should be minimized.
+To enhance the transferability of the pretrained unimanual policy in bimanual manipulation tasks, the observation discrepancy between unimanual and bimanual systems should be minimized.
 We propose a voxel aligner to generate spatial soft masks to highlight relevant visual clues for different arms, whose goal is to align the visual input of the unimanual policy model for each arm with those during the pretraining stage.
@@
 The contributions are summarized as follows:
-  • We propose a model-agnostic plug-and-play framework named AnyBimanual that transfers an arbitrary pre-trained unimanual policy to generalizable bimanual manipulation policy with limited bimanual demonstrations.
+  • We propose a model-agnostic plug-and-play framework named AnyBimanual that transfers an arbitrary pretrained unimanual policy to generalizable bimanual manipulation policy with limited bimanual demonstrations.
   • We introduce a skill manager to dynamically schedule skill representations for unimanual policy transferring and a visual aligner to mitigate the observation discrepancy between unimanual and bimanual systems for transferability enhancement.
@@
 Bimanual manipulation agents [6, 38, 22, 71, 68, 27, 45, 24, 31, 11] are able to handle a large variety of tasks by predicting a trajectory of bimanual operation, which is of great significance in complex applications from household service [72], robotic surgery [33], to component assembly in factories [9].
-To achieve multi-task learning for generalizable Bimanual manipulation, earlier studies attempted to leverage the emerged general understanding and reasoning capacities of pre-trained foundation models like LLMs [53] and VLMs [10], where the foundation model was prompted to generate a high-level plan for low-level executors.
+To achieve multi-task learning for generalizable Bimanual manipulation, earlier studies attempted to leverage the emerged general understanding and reasoning capacities of pretrained foundation models like LLMs [53] and VLMs [10], where the foundation model was prompted to generate a high-level plan for low-level executors.
 However, the performance of directly leveraging foundation models in a training-free manner is bottlenecked by the predefined low-level executor, which struggles to generalize to more contact-rich tasks like straightening a rope.
-To overcome this challenge, robotic foundation models [41, 46, 17, 8, 7, 51] that pre-trained on large-scale real-world demonstrations were proposed under the unimanual setting, which has shown high generalizability across everyday manipulation tasks.
+To overcome this challenge, robotic foundation models [41, 46, 17, 8, 7, 51] that pretrained on large-scale real-world demonstrations were proposed under the unimanual setting, which has shown high generalizability across everyday manipulation tasks.
 However, bimanual tasks demand precise coordination of two high degree-of-freedom arms, making the teleoperation of demonstrations for training generalizable policies also costly.
@@
 While the predefined atomic movements did boost the success rate on specific tasks, they are often difficult for even human users to specify, which restricts the deployment scenarios of these methods.
-In this paper, we propose leveraging learnable skill primitives to represent the learned commonsense of the pre-trained unimanual policy, so that the knowledge can be transferred across different levels of tasks.
+In this paper, we propose leveraging learnable skill primitives to represent the learned commonsense of the pretrained unimanual policy, so that the knowledge can be transferred across different levels of tasks.
 3 Approach
@@
 However, due to the high cost of data collection in bimanual systems, the scarcity of expert demonstrations limits the generalizability of these methods across tasks.
-To address this, we present to transfer pre-trained generalizable unimanual policy for general bimanual manipulation.
+To address this, we present to transfer pretrained generalizable unimanual policy for general bimanual manipulation.
 3.2 Overall Pipeline
 Overall pipeline of our AnyBimanual method is shown in Figure 2.
-For the language branch, we employed a pre-trained text encoder [49] to parse the bimanual instruction to language embeddings with high-level semantics, where the skill manager schedules the skill primitives with composition and compensation to enhance the language embeddings that instruct relevant subtasks for different arms.
+For the language branch, we employed a pretrained text encoder [49] to parse the bimanual instruction to language embeddings with high-level semantics, where the skill manager schedules the skill primitives with composition and compensation to enhance the language embeddings that instruct relevant subtasks for different arms.
 Therefore, the pre-trained unimanual policy model can be prompted to generate feasible manipulation policy for each arm with high generalization ability across tasks with the sharable manipulation knowledge.
@@
 The visual aligner generates a soft spatial mask to align the visual representation of unimanual policy model with its representation during pretraining, so that the observation discrepancy between unimanual and bimanual systems can be minimized for policy transferability enhancement.
-We employ two pre-trained unimanual models to predict the left and right robot actions based on the text embeddings and visual representations, where the pre-trained unimanual policy can be multi-modal transformer-based policy [50, 41, 8, 7, 17] or diffusion-based policy [39, 51].
+We employ two pretrained unimanual models to predict the left and right robot actions based on the text embeddings and visual representations, where the pretrained unimanual policy can be multi-modal transformer-based policy [50, 41, 8, 7, 17] or diffusion-based policy [39, 51].
 3.3 Scheduling Unimanual Skill Primitives
@@
 We start with a discrete primitive skill set $\mathcal{Z}=\{z_{1},z_{2},...,z_{K}\}$, where $K$ is a hyper-parameter that indicates the number of skill primitives.
-To realize end-to-end skill discovery and scheduling, each potential skill is an implicit embedding $z_{k}\in\mathbb{R}^{D}$, which can be initialized with the corresponding language template tokens of the pre-trained unimanual policy to mitigate the domain gap.
+To realize end-to-end skill discovery and scheduling, each potential skill is an implicit embedding $z_{k}\in\mathbb{R}^{D}$, which can be initialized with the corresponding language template tokens of the pretrained unimanual policy to mitigate the domain gap.
 Figure 3: Shareable skills across unimanual and bimanual settings.
@@
 As depicted in Figure 3, considering a bimanual task Handover, it can be explicitly solved by scheduling two unimanual primitive skills, i.e., the left arm Place the block to the right gripper while the right arm Pick it from the left gripper.
-Though every language embedding that passed through the pre-trained single policy can be represented as a linear combination of the skill set, the combination weights that specify the importance of each skill are dynamic across the task completion process.
+Though every language embedding that passed through the pretrained single policy can be represented as a linear combination of the skill set, the combination weights that specify the importance of each skill are dynamic across the task completion process.
 We propose to parametrize a multi-model transformer named skill manager to dynamically predict the combination weight for each arm at each time step.
@@
 Learning Aligned Visual Representations.
-Our goal is to mitigate the visual domain gap between the unimanual and bimanual setting, so that the pre-trained commonsense knowledge in unimanual policy can be transferred with high adaptation ability.
+Our goal is to mitigate the visual domain gap between the unimanual and bimanual setting, so that the pretrained commonsense knowledge in unimanual policy can be transferred with high adaptation ability.
 Since we can not access the unimanual pretraining data in common usages, we instead impose a mutually exclusive prior to the visual aligner.
@@
 3.5 Learning Objectives
-The decomposed skill and volumetric representation are used to pass through the two pre-trained unimanual policies to predict the optimal actions of the two end-effectors.
-We assume access to a pre-trained unimanual policy $p$, which is fundamentally a multi-model multi-task neural network that takes visual and language embedding as inputs and outputs end-effector actions.
-Our AnyBimanual is a model-agnostic plug-and-play method, which indicates that the architecture of pre-trained unimanual policy $p$ is flexible in different architectures such as multi-modal transformer-based policies [50, 41] and diffusion policies [39, 51].
+The decomposed skill and volumetric representation are used to pass through the two pretrained unimanual policies to predict the optimal actions of the two end-effectors.
+We assume access to a pretrained unimanual policy $p$, which is fundamentally a multi-model multi-task neural network that takes visual and language embedding as inputs and outputs end-effector actions.
+Our AnyBimanual is a model-agnostic plug-and-play method, which indicates that the architecture of pretrained unimanual policy $p$ is flexible in different architectures such as multi-modal transformer-based policies [50, 41] and diffusion policies [39, 51].
 To supervise the model with the provided expert demonstrations for behavior cloning, we leverage the cross-entropy loss to learn accurate action prediction for each arm:
@@
 Table 1: Multi-Task Test Results in Simulator.
-Average success rates (%) of general bimanual manipulation agents trained with $20$ or $100$ demonstrations per task and evaluated over $25$ episodes.
+Average success rates (%) of general bimanual manipulation agents trained with $20$ or $100$ demonstrations per task and evaluated over $100$ episodes.
 ‘LF’ means the leader-follower architecture that transfers unimanual manipulation policy for bimanual manipulation.
@@
 […]
-RVT [28]-LF 0 0 0 0 0 0 4 12 0 4 12 16
-RVT-LF + AnyBimanual 0 0 4 8 0 8 4 16 0 8 20 24
-PerAct [50]-LF 0 0 0 4 4 8 4 12 0 4 8 16
-PerAct-LF + AnyBimanual 0 4 0 4 4 4 8 12 0 8 12 24
-PerAct2 [31] 0 8 4 16 0 4 8 40 0 4 0 0
-PerAct2 + Pretraining 0 4 0 20 4 12 16 12 0 12 0 8
-PerAct + AnyBimanual 8 8 12 32 8 24 4 32 4 12 28 16
+RVT [28]-LF 1 2 1 1 1 2 3 3 4 6 7 18
+RVT-LF + AnyBimanual 1 3 2 4 2 5 3 4 4 6 11 21
+PerAct [50]-LF 1 2 3 4 5 11 4 7 6 12 7 17
+PerAct-LF + AnyBimanual 2 2 3 5 12 14 8 8 15 17 23 29
+PerAct2 [31] 3 4 2 4 6 8 4 4 4 3 5 6
+PerAct2 + Pretraining 3 5 5 7 13 17 9 14 2 5 23 31
+PerAct + AnyBimanual 4 7 6 8 17 24 22 36 9 14 31 46
 put in
@@
 […]
-RVT [28]-LF 0 0 12 12 0 0 0 0 0 4 0 0
-RVT-LF + AnyBimanual 0 0 16 20 0 4 0 4 4 4 0 8
-PerAct [50]-LF 0 0 0 8 0 0 8 28 0 0 0 4
-PerAct-LF + AnyBimanual 0 4 4 8 0 0 16 32 0 0 0 8
-PerAct2 [31] 0 0 4 12 8 0 8 0 0 0 28 24
-PerAct2 + Pretraining 0 0 44 64 0 4 0 8 8 4 16 28
-PerAct + AnyBimanual 0 8 48 84 8 8 4 4 8 4 24 28
+RVT [28]-LF 0 0 10 22 0 0 0 0 2 2 3 3
+RVT-LF + AnyBimanual 1 3 17 32 1 2 4 13 1 2 2 7
+PerAct [50]-LF 2 7 2 9 0 3 22 47 0 3 2 5
+PerAct-LF + AnyBimanual 4 9 12 14 1 5 34 57 1 4 3 13
+PerAct2 [31] 7 16 23 41 3 6 25 52 1 3 22 29
+PerAct2 + Pretraining 10 22 27 40 3 7 29 55 3 8 24 33
+PerAct + AnyBimanual 13 26 39 73 7 15 43 67 9 24 31 44
 Table 2: Comparison of AnyBimanual with Different Techniques.
@@
 Row ID Skill Manager Visual Aligner Long Generalized Sync Average
-1 - - 6.86 6.00 12.00 9.00
-2 10.85 32.00 8.00 14.33
-3 12.57 44.00 14.00 18.67
-4 13.71 42.00 13.00 19.00
-5 14.28 46.00 21.00 21.67
+1 - - 16.29 23.50 3.50 14.67
+2 19.57 25.50 9.50 16.75
+3 21.57 44.00 15.50 19.75
+4 23.71 42.00 17.00 25.67
+5 27.29 44.00 25.00 32.00
 4 Experiments
 In this section, we first introduce the experimental setups (Section 4.1).
-Then, we transfer various unimanual methods to bimanual manipulation via AnyBimanual, and compare them with the state-of-the-art approaches to show the superiority (Section 4.2).
-We conduct an ablation study to evaluate the validity of the proposed components (Section 4.3).
-We further interpret the learned skill representation and decomposed volumetric representation by visualization (Section 4.4).
-Finally, we report the real-robot results to demonstrate the effectiveness of AnyBimanual in real-world applications (Section 4.5).
+Then, we transfer various unimanual methods to bimanual manipulation via AnyBimanual, and compare them with the state-of-the-art to show uperiority (Section 4.2).
+We conduct an ablation study to evaluate validity of the proposed components (Section 4.3).
+We further interpret learned skill representation and decomposed volumetric representation by visualization (Section 4.4).
+Finally, we report real-robot results to demonstrate effectiveness of AnyBimanual in real-world applications (Section 4.5).
 4.1 Experiment Setup
@@
 For observation, we employ six cameras with a resolution of $256\times 256$ to cover the entire workspace.
-During the training phase, we provide $20$ or $100$ demonstrations for each task, and we evaluate $25$ episodes per task in the testing set to mitigate the randomness.
+During the training phase, we provide $20$ or $100$ demonstrations for each task, and we evaluate $100$ episodes per task in the testing set to mitigate the randomness.
 Real Robot.
@@
 Table 1 presents a comparison of the average success rates for each task and the average performance is shown in Figure 1.
-Our method achieves the highest overall performance, with an average success rate of $21.67$%, setting a new state-of-the-art in general bimanual manipulation.
+Our method achieves the highest overall performance, with an average success rate of $32.00$%, setting a new state-of-the-art in general bimanual manipulation.
 AnyBimanual leverages the knowledge distilled from unimanual models and successfully transfers it to guide general bimanual manipulation.
-As a result, our method outperforms the cutting-edge bimanual method PerAct2 by a significant improvement of $12.67$% on average.
-With a limited set of $20$ demonstrations, our method still defeats the baseline with a sizable margin of $8.00$%.
+As a result, our method outperforms the cutting-edge bimanual method PerAct2 by a significant improvement of $17.33$% on average.
+With a limited set of $20$ demonstrations, our method still defeats the baseline with a sizable margin of $10.50$%.
 Note that we aim to train multi-task agents to handle a wide range of tasks within a single model, hence the performance is lower than the reported in the original paper.
-Especially, we observe that AnyBimanual shows greater improvement in long-horizon tasks like put in fridge, multi-variations tasks like press buttons, and tasks demand synchronous coordinations like pick plate.
-Moreover, AnyBimanual enables plug-and-play transferring of PerAct-LF and RVT, which also leads to relative boosts of $25.87$% ($7.00$% to $9.00$%) and $116.65$% ($4.00$% to $8.67$%) on average.
+Especially, we observe that AnyBimanual shows greater improvement in long-horizon tasks like put in fridge, multi-variations tasks like press buttons, and tasks demand synchronous coordinations like straighten rope.
+Moreover, AnyBimanual enables plug-and-play transferring of PerAct-LF and RVT, which also leads to relative boosts of $72.76$% ($4.92$% to $8.50$%) and $39.41$% ($10.58$% to $14.75$%) on average.
 However, incorporating AnyBimanual slightly affects the performance on the simple, short-horizon Lift ball, due to the additional complexity in input processing.
@@
 Skill Manager.
-By employing the skill manager in the experiment, we observe that the average success rate increases by $4.67$% (Row $4$ vs.
+By employing the skill manager in the experiment, we observe that the average success rate increases by $8.92$% (Row $4$ vs.
 Row $2$).
@@
 Visual Aligner.
-Additionally, we incorporate the spatial soft masking from the visual aligner, resulting in a substantial performance improvement of 4.34% (Row $3$ vs.
+Additionally, we incorporate the spatial soft masking from the visual aligner, resulting in a substantial performance improvement of 3.00% (Row $3$ vs.
 Row $2$).
@@
 The results demonstrate the visual aligner preserves the generalizability stored in the unimanual policy perfectly, facilitating seamless coordination between the arms.
-By integrating all techniques in our AnyBimanual, the success rate improves from $9.00$% to $21.67$% (Row $5$), validating the importance of properly mining unimanual knowledge for bimanual manipulation.
+By integrating all techniques in our AnyBimanual, the success rate improves from $14.67$% to $32.00$% (Row $5$), validating the importance of properly mining unimanual knowledge for bimanual manipulation.
 4.4 Qualitative Analysis
@@
 5 Conclusion
-In this paper, we have introduced AnyBimanual, a framework designed to transfer pre-trained unimanual manipulation policies to multi-task bimanual manipulation with few bimanual demonstrations.
+In this paper, we have introduced AnyBimanual, a framework designed to transfer pretrained unimanual manipulation policies to multi-task bimanual manipulation with few bimanual demonstrations.
 We develop a skill manager to dynamically schedule skill primitives discovered from unimanual policies, enabling their effective adaptation for bimanual tasks.
@@
 We fuse voxel embeddings and language embeddings by projecting both to a hidden size of $256$ and concatenating them.
-After layer normalization, the combined sequence is processed by a frozen GPT-2 Transformer configured with a hidden size of $128$ and $4$ attention heads.
+After layer normalization, the combined sequence is processed by a frozen GPT-2 Transformer configured with a hidden size of $256$ and $8$ attention heads.
 The transformer’s output is passed through a linear layer to produce logits over 18 classes, and softmaxed to obtain probabilities.
@@
 After element-wise multiplication of these masks with the original input, we incorporate residual connections by adding the masked outputs back to the input embeddings.
+7.3 Skill Primitives
+The skill primitives use CLIP embeddings of unimanual language templates (e.g., ‘open the   drawer’) from Table 1 of the supplementary file, as we expect them to represent foundational motions that generalize across object categories in unseen scenarios.
 8 Additional Experimental Results
@@
 In Figure 12, we cluster the skill representations while evaluating the last checkpoint of PerAct+AnyBimanual in RLBench2.
-We can observe that the learned skill representations are reasonably clustered into $18$ categories, which corresponds to the number of pre-trained unimanual tasks.
+We can observe that the learned skill representations are reasonably clustered into $18$ categories, which corresponds to the number of pretrained unimanual tasks.
 We also visualize the related unimanual skill of each cluster by finding the nearest unimanual task embedding.
@@
 In contrast, our AnyBimanual agent ensures both arms make contact with the box and successfully push it into the red area, demonstrating the superior ability of our method in having each arm correctly identify contact with the object and successfully complete the task.
+8.3 Visualization
+Skill Manager.
+We visualize the skill manager in two challenging tasks requiring heavy coordination.
+(Figure 15) AnyBimanual completes these tasks by scheduling skill representations across arms simultaneously and dynamically.
+Figure 15: Visualization of Skill Manager.
+Viusal Aligner. By predicting spatial soft masks for voxel observation, the visual aligner encourages both arms to focus on the interaction and ignore irrelevant parts. (Figure 16)
 8.4 Error Modes
@@
 We hope that the analysis can enlighten future works in bimanual policy transferring and prevent potential risks in real-world deployments.
+Figure 16: Visualization of Visual Aligner.
+8.5 Bimanual Decomposability
+We complement a heuristic measurement to assess the feasibility of decomposing a given bimanual task with a learned skill manager at the semantic level.
+Specifically, higher entropy in the predicted combination weights indicates greater difficulty in decomposition with certain unimanual primitives.
+In Figure 17, this metric strongly correlates with final success rates in four simulation tasks, delineating the boundary.
+Figure 17: Bimanual Decomposability.
 8.6 Real-world Videos